Benjamin is throwing balls of paper into his trash can. He gets 60% of them into the can. What is the most likely result of his next two attempts?
Answer by stanbon(75887)

Benjamin is throwing balls of paper into his trash can. He gets 60% of them into the can. What is the most likely result of his next two attempts?
---------
make 2: 0.6^2 = 0.36
miss 2: 0.4^2 = 0.16
-----
make 1 : 1 - (0.36+0.16) = 1 - 0.52 = 0.48
--------
Most likely is make 1
==========================
Cheers,
Stan H.